OpenStudy (countrygirl1431):

Solve the following equation, answer as a reduced, mixed number. Then place the correct number in the box provided. 7(x -2) + 3(x + 2) = 5(x - 6)

OpenStudy (anonymous):

You multiply 7 with x and -2, and 3 with x and 2, and 5 with x and -6. Then you get 7x-14+3x+6 = 5x-30, now you just isolate x

OpenStudy (kc_kennylau):

Please expand the brackets first :) 7(x-2) is actually 7x-14 3(x+2) is actually 3x+6 5(x-6) is actually 5x-30 So we would have (7x-14) + (3x+6) = (5x-30) Do you understand this step?

OpenStudy (countrygirl1431):

yes

OpenStudy (madhu.mukherjee.946):

did you do the sum?

OpenStudy (countrygirl1431):

@kc_kennylau

OpenStudy (kc_kennylau):

Then we can expand it furthermore: (7x-14) + (3x+6) is actually (7x-14+3x+6) Therefore (7x-14+3x+6)=(5x-30)

OpenStudy (kc_kennylau):

Now we can group the terms with "x" on the Left Hand Side, and the terms without on the Right Hand Side: 7x+3x-5x = -30+14-6

OpenStudy (madhu.mukherjee.946):

(7x-14) + (3x+6) = (5x-30) 10x-8=5x-30 5x=22 x=22/5

OpenStudy (countrygirl1431):

thanks